OpenStudy (anonymous):

Your club is selling tickets for a game. On Monday the club sold 3 adult tickets and 1 child ticket for a total of $38. On Tuesday the club sold 3 adult tickets and 2 child tickets for a total of $52. Find the price of a child ticket and an adult ticket.

OpenStudy (blurbendy):

you have to set up 2 equations. the first one would be 3a + 1c = 38 can you come up with the equation for the second part?

OpenStudy (anonymous):

yes

OpenStudy (blurbendy):

okay, what do you get

OpenStudy (anonymous):

3a+2c=52

OpenStudy (blurbendy):

okay, so you have: 3a + 1c = 38 3a + 2c = 52 you can use the elimination method or substitution, your choice

OpenStudy (anonymous):

im not real good at those

OpenStudy (blurbendy):

just pick one

OpenStudy (anonymous):

elimination

OpenStudy (blurbendy):

okay, so if we use the elimination method, we can eliminate 3a pretty easily from both equations by multiplying the entire first one by -1

OpenStudy (blurbendy):

what would that give you?

OpenStudy (anonymous):

-3a-1c=-38

OpenStudy (blurbendy):

good, so now you have: -3a-1c=-38 3a + 2c = 52 now add the equations together: -3a-1c=-38 + 3a + 2c = 52

OpenStudy (blurbendy):

what do you get?

OpenStudy (anonymous):

2c=14

OpenStudy (anonymous):

1c

OpenStudy (anonymous):

sry

OpenStudy (blurbendy):

yes, c = 14

OpenStudy (blurbendy):

so now that we know, c = 14, substitute c = 14 back in for c in the top equation

OpenStudy (anonymous):

kk

OpenStudy (blurbendy):

3a + 1(14) = 38

OpenStudy (anonymous):

3a=26

OpenStudy (blurbendy):

38 - 14 = 24

OpenStudy (anonymous):

3a=8

OpenStudy (blurbendy):

3a = 24 a = 8

OpenStudy (anonymous):

lol thats what i meant :)

OpenStudy (blurbendy):

ok good. we found a and c, we're done

OpenStudy (anonymous):

thx

OpenStudy (blurbendy):

yw

OpenStudy (anonymous):

i got more